OpenStudy (luigi0210):

Find the derivative of:

OpenStudy (luigi0210):

\[y= \sqrt {x^3-2x^2}\]

OpenStudy (callisto):

Let \(u = x^3 - 2x^2\) \[\frac{dy}{dx}=\frac{dy}{du}\times \frac{du}{dx}\]\[=\frac{d}{du}(\sqrt{u})\times \frac{d}{dx}(x^2-2x^2)\]Finally, change everything in u back into x.